# Synthetic bootstrapped pretraining

### v1

@@ 3.1 Data, model, and evaluation
To assess the pretraining capability of LM, we measure pretraining test loss and general world knowledge benchmarks. We evaluate held-out test perplexity (exponential of negative log-probability) on 1) OpenWebText2 from EleutherAI (Radford et al., 2018b) and 2) Narrative understanding with LAMBADA (Paperno et al., 2016). We evaluate QA accuracy on 3) Hard scientific reasoning with ARC‑Challenge (Clark et al., 2018); 4) Easy scientific reasoning with ARC‑Easy (Clark et al., 2018); 5) Scientific QA with SciQ (Welbl et al., 2017); 6) Common sense reasoning with Winogrande (Sakaguchi et al., 2021); 7) Reading comprehension with TriviaQA (Joshi et al., 2017); 8) Openbook QA with WebQS (Berant et al., 2013); 9) Broad domain multiple-choice with MMLU (Hendrycks et al., 2020). We directly evaluate the pretrained model with either zero-shot or few-shot prompts. Note that those benchmarks are known to improve significantly with instruction finetuning (Wei et al., 2022). However, we stick to our data-constrained setup and do not introduce any additional data that may confound the comparison.
@@ 4 Experiment results
Table 1: Computed-matched comparison of Synthetic Bootstrapped Pretraining (SBP) and oracle performance gains over the repetition baseline. On average, SBP delivers roughly 47% of the performance improvement in QA accuracy attainable by an oracle with access to 20x more unique data.
@@
MMLU (5-shot) 60.48 +1.26 +2.40 63.93 +0.14 +0.10
Average QA accuracy 49.49 +2.17 +5.09 56.57 +0.74 +1.50
@@ 4.1 Main benchmark performance
At the 200B-scale, we start with the source dataset of $\|{{\mathcal{D}}_{\text{pretrain}}}\|=$10B and curate a SBP dataset of $\|{{\mathcal{S}}_{\text{pretrain}}}\|=$75B tokens (detailed ablation in §A.2). We perform joint training on $\{{{\mathcal{D}}_{\text{pretrain}}},{{\mathcal{S}}_{\text{pretrain}}}\}$ with the principle that we do not repeat any synthetic documents during training. This means that out of a 200B token training budget, we spent 37.5% of it on the 75B synthetic tokens from ${\mathcal{S}}_{\text{pretrain}}$} without any repetition, and the remaining 62.5% on the real dataset ${\mathcal{D}}_{\text{pretrain}}$} repeated 12.5 times. As shown in Table 1, SBP consistently decreases test loss and improves QA accuracy. On average, SBP captures 2.17/5.09=2.17/5.09=42% of the improvement in QA accuracy delivered by the oracle run with 20x additional data access.
@@
Lastly, to validate the benefit of SBP across different training scales, we implement a larger experiment with $\|{{\mathcal{D}}_{\text{pretrain}}}\|=$50B unique tokens under a compute budget of 1T total training tokens. We generate 125B total synthetic tokens $\|{{\mathcal{S}}_{\text{pretrain}}}\|=$125B, with similar ablation presented in §A.2, and adhere to the same no-repetition-for-synthetic-data principle adopted at the 200B-scale. Examining the oracle improvement from Table 1, we can see that the perplexity-based measurements and most QA benchmarks have plateaued at this scale (Liu et al., 2023). However, ARC-Challenge and Winogrande continue to deliver smooth performance changes, making them suitable candidates for tracking pretraining capability at large scales. In particular, in ARC-Challenge, both SBP and Oracle yield a larger performance improvement than their 200B-scale counterparts. That said, other benchmarks still provide a directional signal of capability improvement. This demonstrates the advantage of having a diverse collection of evaluation benchmarks covering a wide range of difficulties. On average, SBP delivers 0.74/1.50=0.74/1.50=49% of the improvement in QA accuracy attained by the oracle.
@@ 7 Acknowledgement
We would like to thank Zonglin Li for guidance on the implementation of ScaNN, Percy Liang for discussion regarding the curation of paired data, and Andrew Ilyas, Bailin Wang, Barry Theobald, Xiang Kong, Yizhe Zhang, Zhiyun Lu for their feedback on the early draft of this paper. ZY was supported by the Albion Walter Hewlett Stanford Graduate Fellowship. EC was supported by the Office of Naval Research grant N00014-24-1-2305, the National Science Foundation grant DMS-2032014, and the Simons Foundation under award 814641. TH was supported by a grant by HAI, DSO labs, gifts from Schmidt Sciences and a grant under the NSF CAREER IIS-2338866, ONR N00014-24-1-2609, and DARPA Cooperative Agreement HR00112520013. This work does not necessarily reflect the position or policy of the government, and no official endorsement should be inferred.

### v2

@@ 3.1 Data, model, and evaluation
To assess the pretraining capability of LM, we measure pretraining test loss and general world knowledge benchmarks. We evaluate held-out test perplexity (exponential of negative log-probability) on 1) OpenWebText2 from EleutherAI (Radford et al., 2018b); 2) Narrative understanding with LAMBADA (Paperno et al., 2016) and 3) Broad domain multiple-choice with MMLU (Hendrycks et al., 2020). We evaluate QA accuracy on 4) Hard scientific reasoning with ARC‑Challenge (Clark et al., 2018); 5) Easy scientific reasoning with ARC‑Easy (Clark et al., 2018); 6) Scientific QA with SciQ (Welbl et al., 2017); 7) Common sense reasoning with Winogrande (Sakaguchi et al., 2021); 8) Reading comprehension with TriviaQA (Joshi et al., 2017); 9) Openbook QA with WebQS (Berant et al., 2013). We directly evaluate the pretrained model with either zero-shot or few-shot prompts. Although MMLU is more commonly known as a QA benchmark, we find that evaluating MMLU accuracy for weak models yields a highly non-smooth readout. As a result, for each MMLU test question, we prepend the question with a 5-shot example of QA pairs and postpend it with the correct answer. Then, we treat each such sample as a text corpus and evaluate LM’s perplexity on such a text sample. Empirically, we find that this perplexity-based MMLU correlates well with MMLU accuracy when the underlying model is large enough to yield a stable readout, and also delivers smooth performance changes for smaller models. Note that those benchmarks are known to improve significantly with instruction finetuning (Wei et al., 2022). However, we stick to our data-constrained setup and do not introduce any additional data that may confound the comparison.
@@ 4 Experiment results
Table 1: Computed-matched comparison of Synthetic Bootstrapped Pretraining (SBP) and oracle performance gains over the repetition baseline. On average, SBP delivers roughly 43% of the performance improvement in QA accuracy attainable by an oracle with access to 20x more unique data.
@@
Average QA accuracy 47.66 +2.32 +5.54 55.35 +0.84 +1.73
@@ 4.1 Main benchmark performance
At the 200B-scale, we start with the source dataset of $\|{{\mathcal{D}}_{\text{pretrain}}}\|=$10B and curate a SBP dataset of $\|{{\mathcal{S}}_{\text{pretrain}}}\|=$75B tokens (detailed ablation in §A.2). We perform joint training on $\{{{\mathcal{D}}_{\text{pretrain}}},{{\mathcal{S}}_{\text{pretrain}}}\}$ with the principle that we do not repeat any synthetic documents during training. This means that out of a 200B token training budget, we spent 37.5% of it on the 75B synthetic tokens from ${\mathcal{S}}_{\text{pretrain}}$} without any repetition, and the remaining 62.5% on the real dataset ${\mathcal{D}}_{\text{pretrain}}$} repeated 12.5 times. As shown in Table 1, SBP consistently decreases test loss and improves QA accuracy. On average, SBP captures 2.32/5.54=2.32/5.54=42% of the improvement in QA accuracy delivered by the oracle run with 20x additional data access.
@@
Lastly, to validate the benefit of SBP across different training scales, we implement a larger experiment with $\|{{\mathcal{D}}_{\text{pretrain}}}\|=$50B unique tokens under a compute budget of 1T total training tokens. We generate 125B total synthetic tokens $\|{{\mathcal{S}}_{\text{pretrain}}}\|=$125B, with similar ablation presented in §A.2, and adhere to the same no-repetition-for-synthetic-data principle adopted at the 200B-scale. Examining the oracle improvement from Table 1, we can see that the perplexity-based measurements and most QA benchmarks have plateaued at this scale (Liu et al., 2023). However, ARC-Challenge and Winogrande continue to deliver smooth performance changes, making them suitable candidates for tracking pretraining capability at large scales. In particular, in ARC-Challenge, both SBP and Oracle yield a larger performance improvement than their 200B-scale counterparts. That said, other benchmarks still provide a directional signal of capability improvement. This demonstrates the advantage of having a diverse collection of evaluation benchmarks covering a wide range of difficulties. On average, SBP delivers 0.84/1.73=0.84/1.73=48% of the improvement in QA accuracy attained by the oracle.
@@ 7 Acknowledgement
We would like to thank Zonglin Li for guidance on the implementation of ScaNN, Percy Liang for discussion regarding the curation of paired data, and Andrew Ilyas, Bailin Wang, Barry Theobald, Xiang Kong, Xin Cheng, Yizhe Zhang, Zhiyun Lu for their feedback on the early draft of this paper. ZY was supported by the Albion Walter Hewlett Stanford Graduate Fellowship. EC was supported by the Office of Naval Research grant N00014-24-1-2305, the National Science Foundation grant DMS-2032014, and the Simons Foundation under award 814641. TH was supported by a grant by HAI, DSO labs, gifts from Schmidt Sciences and a grant under the NSF CAREER IIS-2338866, ONR N00014-24-1-2609, and DARPA Cooperative Agreement HR00112520013. This work does not necessarily reflect the position or policy of the government, and no official endorsement should be inferred.